Find the values for m and b in the following system so that the solution to the system is (-3,4)
5x+7y=b
mx+y=22


The value for m =>


The value for b =>

Answer by ankor@dixie-net.com(22740)

Find values for m and b in the following system so that the solution to the system is (-3,4)
:
The value for b =>
5x + 7y = b
:
Find b by using x,y coordinates given -3, 4
5(-3) + 7(4) = b
-15 + 28 = b
b = + 13
:
The value for m =>
mx + y = 22
Find the slope (m) using the given coordinates x,y; -3,4
m(-3) + 4 = 22
-3m = 22 - 4
-3m = 18
m = 18/-3
m = -6;
: